In 6 years, Nina’s age will be double what it was 6 years ago. Find Nina’s present age.
Answer by ikleyn(52788)

.
N + 6 = 2*(N-6)

N + 6 = 2N - 12

12 + 6 = 2N - N

N = 18.


Answer.  Nina is 18 years old.